OpenStudy (goformit100):

© In a small sweet shop people usually buy either one cake or one box of chocolate. One day shop sold 57 cakes and 36 boxes of chocolates. How many customers were there that day if 12 people bought both a cake and a box of chocolates ?

OpenStudy (goformit100):

@mary.rojas

OpenStudy (mary.rojas):

lol goformit...I dont know if it is right but there: 57-12= 45 36-12= 24 12+45+24=81 81 people